OpenStudy (anonymous):

How would I turn this into a standard ellipses equation? y^2 + 2x^2 = 1

OpenStudy (tkhunny):

It almost is! \(y^{2} + 2x^{2} = 1 \longrightarrow y^{2} + \dfrac{x^{2}}{1/2} = 1 \longrightarrow y^{2} + \dfrac{x^{2}}{\left(1/\sqrt{2}\right)^{2}} = 1\) That's about it.